What is the angle between I+j and i-j
Please answer with full explanation please send me answer as soon as possible

Grade:11

1 Answers

sarthak saxena
5 years ago
| i+j | = | i-j | = 21/2
For Sacalar Product,
(i+j).(i-j) = 21/2x21/2xcost
1-1        = 2.cost
:. cost = o
:. t = 90